# Poly Kernel Inception Network for Remote Sensing Detection

### 10 Mar 2024 v1

@@ 2 Related Work
OBB for Remote Sensing Object Detection. To address the challenge of arbitrary orientations of the objects in RSIs, one research direction focuses on developing specialized OBB detectors. This includes introducing feature refinement techniques into the detector neck [69, 71], extracting the rotated region of interest (RoI) [10, 65], designing specific detection heads for the OBBs [26, 21, 48].etc.
Though improving over general horizontal bounding box (HBB) detectors, these methods often suffer from issues like boundary discontinuity due to their relatively inflexible object representations obtained by augmenting horizontal object representations with additional angle parameters. To mitigate the aforementioned issues, another line of research has been dedicated to developing new object representations for detecting OBBs [67, 31, 62, 76, 15, 70]. For example, Xu et al[67] propose to describe a multi-oriented object by adding four gliding offset variables to classical HBB representation. Li et al[31] characterize oriented objects using a set of points to achieve more accurate orientation estimation. Some others [27, 72, 73, 4] utilize Gaussian distributions to model the OBBs for object detection and design new loss functions [51] to guide the learning process.
@@ 4.1 Experimental Setup
Training. Our training process contains ImageNet [9] pretrain and remote sensing object detector training. For ImageNet pretrain, our PKINet is trained on the ImageNet-1K under the MMPretrain [6] toolbox. In the main experiment, we train it for 300 epochs for higher performance like previous works [65, 71, 50, 32]. In the process of pretrain, we adapt the AdamW [29] optimizer with a momentum of 0.9 and a weight decay of 0.05. Cosine schedule [45] and warm-up strategy are employed to adjust the learning rate. We use 8 GPUs with a batch size of 1024 for pertaining. For remote sensing object detector training, experiments are conducted on MMRotate [84] framework. To compare with other methods, we use trainval sets of these benchmarks and their test sets for testing. Following the settings of previous methods [21, 65, 71, 78], we crop the original images into $1024\times 1024$ patches with overlaps of 200 for DOTA-v1.0 and DOTA-v1.5 datasets. For HRSC2016 and DIOR-R datasets, the input size is set as $800\times 800$. Models are trained with 30 epochs, 30 epochs, 60 epochs, and 36 epochs for DOTA-v1.0, DOTA-v1.5, HRSC2016, and DIOR-R. We employ AdamW [29] optimizer with a weight decay of 0.05. The initial learning rate is set to 0.0002. All the flops reported are calculated when the input image size is $1024\times 1024$. To prevent over-fitting, images undergo random resizing and flipping during the training following previous methods [21, 65, 71, 78]. Five-run mAP of our method are reported for HRSC2016 and DIOR-R.
@@
Reproducibility. Our algorithm is implemented in PyTorch. We use eight NVIDIA RTX 4090 GPUs for ImageNet pretraining and four NVIDIA Tesla V100 GPUs for downstream training and testing. To guarantee the reproducibility of our algorithm, our implementation is available at PKINet.
@@ 4.2 Quantitative Results
Figure 3: Visual results on DOTA [64]. Top: LSKNet [32]; Bottom: our PKINet. See §4.3 for details.
Performance on DOTA-v1.5 [64]. As shown in Table 4, our approach achieves outstanding performance on the more challenging dataset DOTA-v1.5 with minuscule instances, evidencing its efficacy and generalization ability to small objects. Our PKINet outperforms the former state-of-the-art methods, achieving an improvement of 1.21%.
@@ 4.3 Qualitative Results
 Method ResNet-18 [22] PVT-T [63] ConNext-N [43] PKINet-S (ours)
@@ 4.4 Diagnostic Experiments
Cross-Stage Partial Structure. Table 7(e) further explores the impact of the Cross-Stage Partial (CSP) structure. Eliminating CSP leads to exponential increases in both parameters and computational costs (by 211% and 159%, respectively). Reducing stage blocks from (4, 12, 20, 4) to (2, 2, 4, 2) allows models without CSP structure to achieve a similar parameter count as the former one, but results in sub-optimal performance due to fewer blocks.
@@ 5 Discussion and Conclusion
In this paper, we propose Poly Kernel Inception Network (PKINet) for remote sensing object detection, which aims at tackling the challenges posed by considerable variations in object scale and contextual diversity in remote sensing images. PKINet employs parallel depth-wise convolution kernels of various sizes to capture dense texture features effectively across different scales. A Context Anchor Attention mechanism is also introduced to capture long-range contextual information further. We experimentally show that PKINet achieves state-of-the-art performance on four famous remote sensing benchmark datasets. Overall, we wish this work to pave the way for future research.
@@
Acknowledgement. This work was supported by the National Natural Science Foundation of China (No. 62102182, 62202227, 62302217, 62306292), Natural Science Foundation of Jiangsu Province (No. BK20220934, BK20220938, BK20220936), China Postdoctoral Science Foundation (No. 2022M721626, 2022M711635), Jiangsu Funding Program for Excellent Postdoctoral Talent (No. 2022ZB267), Fundamental Research Funds for the Central Universities (No. 30923010303).

### 20 Mar 2024 v2

@@ 2 Related Work
OBB for Remote Sensing Object Detection. To address the challenge of arbitrary orientations of the objects in RSIs, one research direction focuses on developing specialized OBB detectors. This includes introducing feature refinement techniques into the detector neck [69, 71], extracting the rotated region of interest (RoI) [10, 65], designing specific detection heads for the OBBs [26, 21, 48].etc. Though improving over general horizontal bounding box (HBB) detectors, these methods often suffer from issues like boundary discontinuity due to their relatively inflexible object representations obtained by augmenting horizontal object representations with additional angle parameters. To mitigate the aforementioned issues, another line of research has been dedicated to developing new object representations for detecting OBBs [67, 31, 62, 76, 15, 70]. For example, Xu et al[67] propose to describe a multi-oriented object by adding four gliding offset variables to classical HBB representation. Li et al[31] characterize oriented objects using a set of points to achieve more accurate orientation estimation. Some others [27, 72, 73, 4] utilize Gaussian distributions to model the OBBs for object detection and design new loss functions [51] to guide the learning process.
@@ 4.1 Experimental Setup
Training. Our training process contains ImageNet [9] pretrain and remote sensing object detector training. For ImageNet pretrain, our PKINet is trained on the ImageNet-1K under the MMPretrain [6] toolbox. In the main experiment, we train it for 300 epochs for higher performance like previous works [65, 71, 50, 32]. In the process of pretrain, we adapt the AdamW [29] optimizer with a momentum of 0.9 and a weight decay of 0.05. Cosine schedule [45] and warm-up strategy are employed to adjust the learning rate. We use 8 GPUs with a batch size of 1024 for pertaining. For remote sensing object detector training, experiments are conducted on MMRotate [84] framework. To compare with other methods, we use trainval sets of these benchmarks and their test sets for testing. Following the settings of previous methods [21, 65, 71, 78], we crop original images into $1024\times 1024$ patches with overlaps of 200 for DOTA-v1.0 and DOTA-v1.5 datasets. For HRSC2016 and DIOR-R datasets, the input size is set as $800\times 800$. Models are trained with 30 epochs, 30 epochs, 60 epochs, and 36 epochs for DOTA-v1.0, DOTA-v1.5, HRSC2016, and DIOR-R. We employ AdamW [29] optimizer with a weight decay of 0.05. The initial learning rate is set to 0.0002. All flops reported are calculated when the input image size is $1024\times 1024$. To prevent over-fitting, images undergo random resizing and flipping during training following previous methods [21, 65, 71, 78]. Five-run average mAP of our method are reported for HRSC2016 and DIOR-R.
@@
Reproducibility. Our algorithm is implemented in PyTorch. We use eight NVIDIA RTX 4090 GPUs for ImageNet pretraining and four NVIDIA Tesla V100 GPUs for downstream training and testing.
@@ 4.2 Quantitative Results
Figure 3: Visual results on DOTA-v1.0 dataset [64]. Top: LSKNet [32]; Bottom: our PKINet. See §4.3 for details.
Performance on DOTA-v1.5 [64]. As shown in Table 4, our approach achieves outstanding performance on the more challenging dataset DOTA-v1.5 with minuscule instances, evidencing its efficacy and generalization ability to small objects. Our PKINet outperforms the former state-of-the-art methods, achieving an improvement of 1.21%. C.
@@ 4.3 Qualitative Results
 Method ResNet-18 [22] PVT-T [63] ConvNeXt-N [43] PKINet-S (ours)
@@ 4.4 Diagnostic Experiments
Cross-Stage Partial Structure. Table 7(e) further explores the impact of the Cross-Stage Partial (CSP) structure. Eliminating CSP leads to exponential increases in both parameters and computational costs (by 211% and 159%, respectively). By reducing the number of blocks in each stage from (4, 12, 20, 4) to (2, 2, 4, 2), models without CSP structure can reach a similar parameter as the former one. However, sub-optimal performance is achieved due to the decreased number of stage blocks.
@@ 5 Discussion and Conclusion
In this paper, we propose Poly Kernel Inception Network (PKINet) for remote sensing object detection, which aims at tackling the challenges posed by considerable variations in object scale and contextual diversity in remote sensing images. PKINet employs parallel depth-wise convolution kernels of various sizes to capture dense texture features effectively across different scales. A Context Anchor Attention mechanism is also introduced to capture long-range contextual information further. We experimentally show that PKINet achieves state-of-the-art performance on four famous remote sensing benchmark datasets.